OpenStudy (anonymous):

You are expecting a call from a friend anytime between 3:00 P.M. and 5:00 P.M. At 3:20 P.M, you discover that someone accidentally left the phone off the hook. What is the probability that you missed your friend’s call?

OpenStudy (anonymous):

No. of minutes between 3 : 00 PM to 5 :00 PM = ?

OpenStudy (anonymous):

i dont think so..

OpenStudy (anonymous):

my answers are... about 10%

OpenStudy (anonymous):

about 33% about 17% about 40%

OpenStudy (anonymous):

P(given event) = \(\cfrac{\textbf{ No. of events in favor} }{ \textbf{Total events} }\) So here : P(given event) = \(\cfrac{\textbf{20 minutes}}{\textbf{total minutes between 3:00 PM to 5:00 PM}}\)

OpenStudy (anonymous):

See ... first of all total minutes between 3:00 PM to 5:00 PM would be 120 minutes, right?

OpenStudy (anonymous):

honestly i coulnt tell you.. i suck at math...

OpenStudy (anonymous):

See there are two hours between 3 PM and 5 PM ... and each hour contain 60 minutes. So total minutes = 120 minutes (from 3 PM to 5 PM)

OpenStudy (anonymous):

Now, So we have : \(\cfrac{20}{120} \) = P(E) \(\cfrac{1}{6} = P(E)\) Now we will calculate the percentage... by multiplying it by 100%

OpenStudy (anonymous):

That is \(\cfrac{1}{6} \times 100 \% = \cfrac{100}{6} \% = P(E)\) You can calculate it now..

OpenStudy (anonymous):

what is the P(E)

OpenStudy (anonymous):

Probability of the given event.

OpenStudy (anonymous):

about 33% it will be..

OpenStudy (anonymous):

ok im tryin to work it out.. but im like lost.. how did you get that ?

OpenStudy (anonymous):

Is it clear for you that it will be \(\cfrac{100}{6} \% \) ?

OpenStudy (anonymous):

yea.. cause you would have to multiply it by 100.. right ?

OpenStudy (anonymous):

Yes! Now calculate : 100/6 Sorry it will be about 17%

OpenStudy (anonymous):

so since 20/120 is 1/6 you would divide 100/6 .. ..

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

Yes ... as 1/6 was multiplied by 100 then.. 100/6 = ?

OpenStudy (anonymous):

16.67 which woul be 50/3

OpenStudy (anonymous):

Yeah! you have it as about 17%

OpenStudy (anonymous):

oooooooo ok i see

OpenStudy (anonymous):

omg thank you..

OpenStudy (anonymous):

You're welcome @ohnaw . Best of luck